# fastkqr: A Fast Algorithm for Kernel Quantile Regression

### v1

@@ Abstract
Quantile regression is a powerful tool for robust and heterogeneous learning that has seen applications in a diverse range of applied areas. However, its broader application is often hindered by the substantial computational demands arising from the non-smooth quantile loss function. In this paper, we introduce a novel algorithm named fastkqr, which significantly advances the computation of quantile regression in reproducing kernel Hilbert spaces. The core of fastkqr is a finite smoothing algorithm that magically produces exact regression quantiles, rather than approximations. To further accelerate the algorithm, we equip fastkqr with a novel spectral technique that carefully reutilizes matrix computations. In addition, we extend fastkqr to accommodate a flexible kernel quantile regression with a data-driven crossing penalty, addressing the interpretability challenges of crossing quantile curves at multiple levels. We have implemented fastkqr in a publicly available R package. Extensive simulations and real applications show that fastkqr matches the accuracy of state-of-the-art algorithms but can operate up to an order of magnitude faster.
@@ 1 Introduction
Despite its popularity, one primary limitation of quantile regression is its high computational cost, which is also inherited from its median regression origins. This computational burden is mainly due to its non-smooth loss function. To address this, linear quantile regression is often formulated as a linear programming program and solved using the simplex method (Koenker and d’Orey, 1994) or the interior point algorithm (Portnoy and Koenker, 1997). However, the computation becomes more challenging when it comes to the kernel quantile regression (KQR, Takeuchi et al., 2006; Li et al., 2007), the method that is essential for estimating non-linear conditional quantile functions. KQR is typically solved using the interior point method, which has been implemented in the state-of-the-art R package kernlab (Karatzoglou et al., 2004), but in principle, the algorithm only provides approximate solutions to the original problem. A seminal work finding the exact solution of linear quantile regression was developed by Chen (2007). However, Chen’s algorithm works only for linear quantile regression, and extending it to KQR is not straightforward. Alternatively, one can consider the least angle regression (LARS) algorithm for computing the exact solution paths of KQR (Hastie et al., 2004; Li et al., 2007; Takeuchi et al., 2009), but it is empirically not as fast as kernlab. Recently, the kernel convolution technique has been used to smooth the quantile regression to efficiently find the approximate solutions (Fernandes et al., 2021; Tan et al., 2022; He et al., 2023).
The first main contribution of this work is the development of a fast algorithm called fastkqr to alleviate the computational burden of KQR. Our core strategy involves smoothing the original problem and recovering the exact solution by leveraging some unique properties of the quantile loss. To solve the smoothed problem, we introduce a novel spectral technique that builds upon the accelerated proximal gradient descent. With this technique, the algorithm operates at a complexity of only $\mathcal{O}(n^{2})$ after an initial eigen-decomposition of the kernel matrix. This efficient implementation makes our algorithm scalable for the KQR computation that involves numerous tuning parameters with different quantile levels.
@@
Figure 1: Fitting quantile curves at various levels using the GAGurine data. In both panels, gray dots represent the data points. The quantile curves are fitted at levels $\tau=0.1,0.3,0.5,0.7$, and $0.9$. The top panel shows quantile curves fitted individually at each level and the gray bands highlight the zones where crossings occur. The bottom panel displays the quantile curves fitted simultaneously using the NCKQR model, where no crossings are present.
Our second major contribution addresses the crossing issue in KQR by enhancing the capabilities of fastkqr to produce non-crossing kernel quantile regression (NCKQR). Following the approach of a non-crossing constraint (Takeuchi et al., 2006; Bondell et al., 2010; Liu and Wu, 2011), we adopt a soft non-crossing penalty. This soft penalty not only eases computation but also adds flexibility, allowing the degree of crossing to be adjusted according to specific scenario tolerances. To manage the increased computational load, we fit NCKQR using a specialized majorization-minimization algorithm featuring two majorization steps. The resulting NCKQR algorithm efficiently tackles the non-crossing penalty, maintaining the same level of computational complexity as fitting a single-level KQR. To give a quick demonstration, the bottom panel of Figure 1 depicts the five quantile curves fitted using NCKQR, showcasing that no crossings occur.
Our numerical studies demonstrate that fastkqr can be significantly faster than the state-of-the-art solver, kernlab. For example, we fit KQR using fastkqr and kernlab on simulated data considered by Yuan (2006). With a sample size $n=1000$, fastkqr completed the computation in 70 seconds. In contrast, kernlab, while maintaining comparable accuracy with fastkqr, took about 700 seconds. We have implemented our algorithms in an R package fastkqr, which is publicly available on the Comprehensive R Archive Network (CRAN) at https://CRAN.R-project.org/package=fastkqr.
@@ 2.1 A Smooth surrogate for kernel quantile regression
and each $\mathbf{K}_{i}^{\top}$ is the $i$th row of $\mathbf{K}$, the $n\times{}n$ symmetric kernel matrix whose $(i,j)$th entry is $K(\mathbf{x}_{i},\mathbf{x}_{j})$.
Solving problem (2) is challenging primarily because the check loss function is non-smooth. To efficiently compute KQR, we propose to first smooth the original problem. Given $\tau\in(0,1)$, we introduce a $\gamma$-smoothed check loss function,
@@
where $\gamma>0$ is the smoothing parameter. We then give the following smooth surrogate of the objective function in problem (2):
@@ 2.2 An exact finite smoothing algorithm
Then, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$ holds, where $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the solution of problem (2).
@@
$\begin{split}&(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha}),\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in{}S.\end{split}$
@@
Theorem 2 establishes that, by the set expansion method, any set $S\subseteq{}{}S_{0}$ will either remain unchanged or be expanded towards the singular set $S_{0}$. Therefore, we can initiate the process with the empty set $\hat{S}=\emptyset\subseteq S_{0}$ and keep expanding $\hat{S}$ by iteratively applying the set expansion method until $\hat{S}$ remains unchanged, i.e., $\hat{S}=\mathcal{E}(\hat{S})$. Since the sample size is finite, $\hat{S}$ becomes unchanged in finite steps. We thus name the whole process the finite smoothing algorithm.
@@
Therefore, with the finite smoothing algorithm, the exact solution of problem (2) can be obtained by iteratively solving problem (5) and augmenting the set $\hat{S}$. Since $\gamma^{*}$ is still unknown in practice, we handle this by repeatedly implementing the above procedure with a decreasing sequence of values of $\gamma$. The algorithm is terminated once a solution satisfies the Karush–Kuhn–Tucker (KKT) conditions of problem (2). In our implementation, we initiate this process with $\gamma=1$ and iteratively update it by reducing $\gamma$ to a quarter of its previous value, i.e., $\gamma\leftarrow\gamma/4$. This approach generally converges after only three or four iterations of updating $\gamma$.
@@ 2.3 Solving the smoothed kernel quantile regression
Let $(b^{(1)},\boldsymbol{\alpha}^{(1)})$ be the initial value of $(b,\boldsymbol{\alpha})$. For each $k=1,2,\ldots$, let $\mathbf{z}^{(k)}$ be the $n$-vector with the $i$th element $H^{\prime}_{\gamma,\tau}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})$ and compute
$\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)$ $=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\frac{1}{4n\gamma}\left\|b\mathbf{1}+\mathbf{K}\boldsymbol{\alpha}-(b^{(k)}\mathbf{1}{}+\mathbf{K}\boldsymbol{\alpha}^{(k)}+{}2\gamma\mathbf{z}^{(k)})\right\|_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}$ (6)
$=\left(\begin{array}[]{c}b^{(k)}\\ \boldsymbol{\alpha}^{(k)}\end{array}\right)+2\gamma\mathbf{P}_{\gamma,\lambda}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}-n\lambda\mathbf{K}\boldsymbol{\alpha}^{(k)}\end{array}\right),$
@@
$\mathbf{P}_{\gamma,\lambda}=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{K}\\ \mathbf{K}^{\top}\mathbf{1}&\mathbf{K}^{\top}\mathbf{K}+2n\gamma\lambda\mathbf{K}\end{array}\right).$
The step size is directly obtained from the following Lipschitz condition of $H_{\gamma,\tau}^{\prime}(t)$,
In fact, the update in equation (6) is essentially equivalent to performing a step of the majorization-minimization (MM) algorithm (Hunter and Lange, 2000),
$\begin{split}\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}&\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})\\ &+\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}^{\prime}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})(b^{(k)}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})\\ &+\frac{1}{4n\gamma}\lVert{}b^{(k)}\mathbf{1}+\mathbf{K}\boldsymbol{\alpha}^{(k)}-b\mathbf{1}-\mathbf{K}\boldsymbol{\alpha}\rVert_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha},\end{split}$
whose objective function majorizes the function $G^{\gamma}(b,\boldsymbol{\alpha})$.
@@ 2.4 A fast spectral technique
To accelerate the APGD algorithm, we develop a spectral technique. We begin with the eigen-decomposition of the kernel matrix, $\mathbf{K}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\boldsymbol{\Lambda}$ is diagonal and $\mathbf{U}$ is orthogonal. Note that this step is free of the parameters $\gamma$ and $\lambda$. Define $\boldsymbol{\Pi}_{\gamma,\lambda}=\boldsymbol{\Lambda}^{2}+2n\gamma\lambda\boldsymbol{\Lambda}$, $\mathbf{v}=\mathbf{U}\boldsymbol{\Lambda}\Pi_{\gamma,\lambda}^{-1}\mathbf{U}^{\top}\mathbf{1}$, and $g=1/\bigl{(}n\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\Pi_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}\bigr{)}$. We have
$\begin{split}\mathbf{P}_{\gamma,\lambda}^{-1}&=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\\ \mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}&\mathbf{U}\Pi_{\gamma,\lambda}\mathbf{U}^{\top}\end{array}\right)^{-1}=g\Biggl{(}\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\Biggr{)}\Bigl{(}\begin{array}[]{ll}1&-\mathbf{v}^{\top}\end{array}\Bigr{)}+\left(\begin{array}[]{cc}0&\mathbf{0}^{\top}\\ \mathbf{0}&\mathbf{U}\Pi_{\lambda}^{-1}\mathbf{U}^{\top}\end{array}\right).\end{split}$ (9)
@@
$\mathbf{P}_{\gamma,\lambda}^{-1}\boldsymbol{\zeta}=g\left\{\mathbf{1}^{\top}\mathbf{z}-\mathbf{v}^{\top}\mathbf{K}\left(\mathbf{z}+n\lambda\boldsymbol{\alpha}\right)\right\}\Biggl{(}\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\Biggr{)}+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}+n\lambda\boldsymbol{\alpha}\right)\end{array}\right),$ (10)
of which the computational complexity is only $\mathcal{O}\left(n^{2}\right)$ when computed from right to left, that is, by only performing matrix-vector multiplications.
In our implementation, we use a warm-start strategy to further amplify the effect of the spectral technique. Specifically, we solve problem (2) with a sequence of tuning parameters $\lambda^{[1]},\lambda^{[2]},\ldots,\lambda^{[L]}$ to obtain the corresponding solutions $(\hat{b}^{[1]},\hat{\boldsymbol{\alpha}}^{[1]}),(\hat{b}^{[2]},\hat{\boldsymbol{\alpha}}^{[2]}),\ldots,(\hat{b}^{[L]},\hat{\boldsymbol{\alpha}}^{[L]})$. With $l>1$, $(\hat{b}^{[l-1]},\hat{\boldsymbol{\alpha}}^{[l-1]})$ is summoned to initialize the finite smoothing and APGD algorithms to solve $(\hat{b}^{[l]},\hat{\boldsymbol{\alpha}}^{[l]})$. Therefore, thanks to both the warm start and the spectral technique, except for the only step of eigen-decomposition, costing the $\mathcal{O}(n^{3})$ complexity, the rest of the entire KQR algorithm to solve problem (2) is only $\mathcal{O}(n^{2})$, hence the speed is significantly enhanced.

### v2

@@ Abstract
Quantile regression is a powerful tool for robust and heterogeneous learning that has seen applications in a diverse range of applied areas. However, its broader application is often hindered by the substantial computational demands arising from the non-smooth quantile loss function. In this paper, we introduce a novel algorithm named fastkqr, which significantly advances the computation of quantile regression in reproducing kernel Hilbert spaces. The core of fastkqr is a finite smoothing algorithm that magically produces exact regression quantiles, rather than approximations. To further accelerate the algorithm, we equip fastkqr with a novel spectral technique that carefully reuses matrix computations. In addition, we extend fastkqr to accommodate a flexible kernel quantile regression with a data-driven crossing penalty, addressing the interpretability challenges of crossing quantile curves at multiple levels. We have implemented fastkqr in a publicly available R package on CRAN. Extensive simulations and real applications show that fastkqr matches the accuracy of state-of-the-art algorithms but can operate up to an order of magnitude faster.
@@ 1 Introduction
Despite its popularity, one primary limitation of quantile regression is its high computational cost, which is also inherited from its median regression origins. This computational burden is mainly due to its non-smooth loss function. To address this, linear quantile regression is often formulated as a linear program and solved using the simplex method (Koenker and d’Orey, 1994) or the interior point algorithm (Portnoy and Koenker, 1997). However, computation becomes more challenging when it comes to kernel quantile regression (KQR, Takeuchi et al., 2006; Li et al., 2007), the method that is essential for estimating non-linear conditional quantile functions. KQR is typically solved using the interior point method, which has been implemented in the state-of-the-art R package kernlab (Karatzoglou et al., 2004), but in principle, the algorithm only provides approximate solutions to the original problem. A seminal work finding the exact solution of linear quantile regression was developed by Chen (2007). However, Chen’s algorithm works only for linear quantile regression, and extending it to KQR is not straightforward. Alternatively, one can consider the least angle regression (LARS) algorithm for computing the exact solution paths of KQR (Hastie et al., 2004; Li et al., 2007; Takeuchi et al., 2009), but it is empirically not as fast as kernlab. Recently, the kernel convolution technique has been used to smooth the quantile regression to efficiently find approximate solutions (Fernandes et al., 2021; Tan et al., 2022; He et al., 2023).
The first main contribution of this work is the development of a fast algorithm called fastkqr to alleviate the computational burden of KQR. Our core strategy involves smoothing the original problem and recovering the exact solution by leveraging some unique properties of the quantile loss. To solve the smoothed problem, we introduce a novel spectral technique that builds upon the accelerated proximal gradient descent algorithm. With this technique, the algorithm operates at a complexity of only $\mathcal{O}(n^{2})$ after an initial eigen-decomposition of the kernel matrix. This efficient implementation makes our algorithm scalable for the KQR computation that involves numerous tuning parameters with different quantile levels.
@@
Figure 1: Fitting quantile curves at various levels using the GAGurine data. In both panels, gray dots represent the data points. The quantile curves are fitted at levels $\tau=0.1,0.3,0.5,0.7$, and 0.90.90.9. The top panel shows quantile curves fitted individually at each level and the gray bands highlight the zones where crossings occur. The bottom panel displays the quantile curves fitted simultaneously using the NCKQR model, where no crossings are present.
Our second major contribution addresses the crossing issue in KQR by enhancing the capabilities of fastkqr to produce non-crossing kernel quantile regression (NCKQR). Following the approach of a non-crossing constraint (Takeuchi et al., 2006; Bondell et al., 2010; Liu and Wu, 2011), we adopt a soft non-crossing penalty. This soft penalty not only eases computation but also adds flexibility, allowing the degree of crossing to be adjusted according to specific scenario tolerances. To manage the increased computational load, we fit NCKQR using a specialized majorization-minimization algorithm featuring two majorization steps. The resulting NCKQR algorithm efficiently tackles the non-crossing penalty, maintaining the same level of computational complexity as fitting a single-level KQR. To give a quick demonstration, the bottom panel of Figure 1 depicts the five quantile curves fitted using NCKQR, showing no crossings.
Our numerical studies demonstrate that fastkqr can be significantly faster than the state-of-the-art solver, kernlab. For example, we fit KQR using fastkqr and kernlab on simulated data considered by Yuan (2006). With a sample size $n=1000$, fastkqr completed the computation in 30 seconds. In contrast, kernlab, while maintaining comparable accuracy with fastkqr, took about 400 seconds. We have implemented our algorithms in an R package fastkqr, which is publicly available on the Comprehensive R Archive Network (CRAN) at https://CRAN.R-project.org/package=fastkqr.
@@ 2.1 A Smooth surrogate for kernel quantile regression
[…]
Solving problem (2) is challenging primarily because the check loss function is non-smooth. To efficiently compute KQR, we propose to first smooth the original problem. Given $\tau\in(0,1)$, we consider a $\gamma$-smoothed check loss function,
@@
where $\gamma>0$ is the smoothing parameter. Similar smoothing strategies for support vector machine have previously been employed in Wang and Zou (2022). A plot of the function $H_{\gamma,\tau}(t)$ is provided in Figure S1 in the online supplement.
We then introduce the following smooth surrogate of the objective function in problem (2):
@@
Note that the above $\gamma$-smoothed function matches the original check loss function in the regions where $t<-\gamma$ and $t>\gamma$, while in the intermediate region $-\gamma\leq t\leq\gamma$, it is smoothed using a quadratic function to ensure continuity. This smoothing approach is natural and simple and thus has been chosen in our implementation. Note that the $\gamma$-smoothed function is not the only option. In Section 2.5, we shall discuss other smoothing approaches.
@@ 2.2 An exact finite smoothing algorithm
Then, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$ holds, where $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the solution to problem (2).
@@
$(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha}),\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in{}S.$
@@
Theorem 2 establishes that, by the set expansion method, any set $S\subseteq{}{}S_{0}$ will either remain unchanged or expand toward the singular set $S_{0}$. Therefore, we can initiate the process with the empty set $\hat{S}=\emptyset\subseteq S_{0}$ and continue expanding $\hat{S}$ by iteratively applying the set expansion method until $\hat{S}$ remains unchanged, i.e., $\hat{S}=\mathcal{E}(\hat{S})$. Since the sample size is finite, $\hat{S}$ becomes unchanged in finite steps. We therefore name the whole procedure the finite smoothing algorithm.
@@
Therefore, with the finite smoothing algorithm, the exact solution of problem (2) can be obtained by iteratively solving problem (5) and augmenting the set $\hat{S}$. Since $\gamma^{*}$ is still unknown in practice, we handle this by repeatedly implementing the above procedure with a decreasing sequence of values of $\gamma$. The algorithm is terminated once a solution satisfies the Karush–Kuhn–Tucker (KKT) conditions of problem (2). In our implementation, we initiate this process with $\gamma=1$ and iteratively update it by reducing $\gamma$ to a quarter of its previous value, i.e., $\gamma\leftarrow\gamma/4$. We observe that this approach typically converges within only three or four iterations of updating $\gamma$.
@@ 2.3 Solving the smoothed kernel quantile regression
First, note that $H_{\gamma,\tau}^{\prime}(t)$ is Lipschitz continuous, that is,
Let $(b^{(1)},\boldsymbol{\alpha}^{(1)})$ be the initial value of $(b,\boldsymbol{\alpha})$. For each $k=1,2,\ldots$, the proximal gradient method updates $(b^{(k+1)},\boldsymbol{\alpha}^{(k+1)})$ by the majorization-minimization principle (Hunter and Lange, 2000),
$\begin{split}\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}&\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})\\ &+\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}^{\prime}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})(b^{(k)}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})\\ &+\frac{1}{4n\gamma}\lVert{}b^{(k)}\mathbf{1}+\mathbf{K}\boldsymbol{\alpha}^{(k)}-b\mathbf{1}-\mathbf{K}\boldsymbol{\alpha}\rVert_{2}^{2}\\ =\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}&\frac{1}{4n\gamma}\left\|\mathbf{K}\boldsymbol{\alpha}+b\mathbf{1}-(\mathbf{K}\boldsymbol{\alpha}^{(k)}+b^{(k)}\mathbf{1}{}+{}\gamma\mathbf{z}^{(k)})\right\|_{2}^{2}+\frac{\lambda}{2}\alpha^{\top}\mathbf{K}\alpha,\end{split}$ (6)
where $\mathbf{z}^{(k)}$ is an $n$-vector whose $i$th element is $H^{\prime}_{\gamma,\tau}(y_{i}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}-b^{(k)})$. Then, we have
$\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)=\left(\begin{array}[]{c}b^{(k)}\\ \boldsymbol{\alpha}^{(k)}\end{array}\right)+\gamma\mathbf{P}_{\gamma,\lambda}^{-1}\boldsymbol{\zeta},$ (7)
@@
$\boldsymbol{\zeta}=\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}-n\lambda\mathbf{K}\boldsymbol{\alpha}^{(k)}\end{array}\right),\mathbf{P}_{\gamma,\lambda}=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{K}\\ \mathbf{K}^{\top}\mathbf{1}&\mathbf{K}^{\top}\mathbf{K}+n\gamma\lambda\mathbf{K}\end{array}\right).$
@@ 2.4 A fast spectral technique
To accelerate the APGD algorithm, we develop a spectral technique, which begins with the eigen-decomposition of the kernel matrix, $\mathbf{K}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\boldsymbol{\Lambda}$ is diagonal and $\mathbf{U}$ is orthogonal. Note that this step is free of the parameters $\gamma$ and $\lambda$. Define $\boldsymbol{\Pi}_{\gamma,\lambda}=\boldsymbol{\Lambda}^{2}+2n\gamma\lambda\boldsymbol{\Lambda}$, $\mathbf{v}=\mathbf{U}\boldsymbol{\Lambda}\Pi_{\gamma,\lambda}^{-1}\mathbf{U}^{\top}\mathbf{1}$, and $g=1/\bigl{(}n\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\Pi_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}\bigr{)}$. Using the Woodbury matrix identity, we obtain
$\begin{split}\mathbf{P}_{\gamma,\lambda}^{-1}&=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\\ \mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}&\mathbf{U}\Pi_{\gamma,\lambda}\mathbf{U}^{\top}\end{array}\right)^{-1}=g\Biggl{(}\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\Biggr{)}\Bigl{(}\begin{array}[]{ll}1&-\mathbf{v}^{\top}\end{array}\Bigr{)}+\left(\begin{array}[]{cc}0&\mathbf{0}^{\top}\\ \mathbf{0}&\mathbf{U}\boldsymbol{\Pi}_{\lambda}^{-1}\mathbf{U}^{\top}\end{array}\right).\end{split}$ (10)
@@
$\mathbf{P}_{\gamma,\lambda}^{-1}\bar{\boldsymbol{\zeta}}=g\left\{\mathbf{1}^{\top}\bar{\mathbf{z}}-\mathbf{v}^{\top}\mathbf{K}\left(\bar{\mathbf{z}}+n\lambda\boldsymbol{\alpha}\right)\right\}\Biggl{(}\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\Biggr{)}+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\bar{\mathbf{z}}+n\lambda\boldsymbol{\alpha}\right)\end{array}\right),$ (11)
where $\bar{\boldsymbol{\zeta}}$ was defined in equation (8). The computational complexity is only $\mathcal{O}\left(n^{2}\right)$ when computed from right to left, that is, by only performing matrix-vector multiplications.
In our implementation, we use a warm-start strategy to further amplify the effect of the spectral technique. Specifically, we solve problem (2) with a sequence of tuning parameters $\lambda^{[1]},\lambda^{[2]},\ldots,\lambda^{[L]}$ to obtain the corresponding solutions $(\hat{b}^{[1]},\hat{\boldsymbol{\alpha}}^{[1]}),(\hat{b}^{[2]},\hat{\boldsymbol{\alpha}}^{[2]}),\ldots,(\hat{b}^{[L]},\hat{\boldsymbol{\alpha}}^{[L]})$. With $l>1$, $(\hat{b}^{[l-1]},\hat{\boldsymbol{\alpha}}^{[l-1]})$ is summoned to initialize the finite smoothing and APGD algorithms to solve for $(\hat{b}^{[l]},\hat{\boldsymbol{\alpha}}^{[l]})$. Therefore, thanks to both the warm start and the spectral technique, except for the only step of eigen-decomposition, which costs an $\mathcal{O}(n^{3})$ complexity, the rest of the entire KQR algorithm to solve problem (2) takes only $\mathcal{O}(n^{2})$, hence the speed is significantly enhanced.
@@
2.5 Alternative smoothing surrogates
Now that we have shown the exact KQR solution to problem (2) can be recovered through a smoothing surrogate, the proposed $\gamma$-smoothed check loss is, however, not the only option that leads to the exact solution. The following theorem provides a broader perspective.
Theorem 5.
Suppose $S_{0}$, the singular set of problem (2), is known. Define
$G^{\gamma}(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha},$
where $H_{\gamma,\tau}(t)$ is a function satisfying the following constraints,
$\begin{split}&1.\enskip{}H_{\gamma,\tau}^{\prime}\left(t\right)\in\partial\rho_{\tau}\left(t\right)\text{ if }t=0,\\ &2.\enskip{}H_{\gamma,\tau}^{\prime}\left(t\right)=\partial\rho_{\tau}\left(t\right)\text{ if }t\neq 0.\end{split}$ (12)
Let
$(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha}),\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in{}S_{0}.$
Then, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$ holds, where $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the solution to problem (2).
Theorem 5 provides a general condition for the smooth surrogate to yield the exact solution to the original non-smooth problem, once the singular set $S_{0}$ is known. Using the same set expansion technique introduced in Section 2.2, we can identify the singular set $S_{0}$ in practice. The general condition can be satisfied by several popular smoothing techniques, including the Moreau envelope (Moreau, 1965; Chen, 2007), Nesterov’s smoothing (Nesterov, 2005), Huber approximation (Yi and Huang, 2017) and kernel density convolution (Tan et al., 2022; He et al., 2023); further details can be found in the supplementary material. Our experiments indicate that their performance is highly comparable, so we adopt the $\gamma$-smoothed check loss in our implementation for simplicity.